OpenStudy (anonymous):

Is 25, 75, 100.... geometric sequence? Please Explain how to get the common ratio.

OpenStudy (anonymous):

To prove if that is a geometric sequence, they must have the same common ratio. So, 100/75 must be equal to 75/25 100/75 = 1.333 while 75/25= 3 Since it is both not equal to each other. The above is not a geometric sequence. Now how to find common ratio. Given a valid geometric sequence like, 1,2,4,8,16. U realize that, \[\frac{16}{8} = \frac{8}{4} = \frac{4}{2}= \frac {2}{1}\] and they all equals to 2. So 2 is the common ratio here.

OpenStudy (anonymous):

So since its not a geometric sequence and it cant be a arithmetic sequence because the numbers arent equal then it is neither right?

OpenStudy (anonymous):

yea it is neither.

OpenStudy (anonymous):

ok thank you very much.

OpenStudy (anonymous):

Np!